a candy store owner wants to mix jelly beans costing $4 per pound with some cherry sours costing $3.50 per pound to obtain a 10-pound mixture costing $3.85 per pound. How many pounds of each type on candy should be used?
Answer by ikleyn(52778)

.
a candy store owner wants to mix jelly beans costing $4 per pound with some cherry sours costing $3.50 per pound
to obtain a 10-pound mixture costing $3.85 per pound. How many pounds of each type on candy should be used?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Solution 1. Two equations

Let x = the amount of the $4 jelly beans and y = the amount of the $3.50 cherry sours (in pounds) to mix.

Then you have two equations 

 x +     y = 10,
4x + 3.50y = 3.85*10.

Simplify:

 x +     y = 10,      (1)
4x + 3.50y = 38.5.    (2)

Express x = 10-y from (1) and substitute it into (2). You will get

4(10-y) + 3.5y = 38.5.

Simplify and solve for y.


Solution 2. One equations

Let x = the amount of the $4 jelly beans (in pounds).
Then the amount of the $3.50 cherry sours (in pounds) is 10-x.

The equation is

4x + 3.50*(10-x) = 3.85*10.

Simplify and solve for x.

There is entire bunch of the lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts
    - Word problems on mixtures for dry substances like candies, dried fruits
    - Word problems on mixtures for dry substances like soil and sand
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Mixture problems".